OpenStudy (anonymous):

sketch the graph of the function over the interval [0,2π], by comparing it to the graph of y=sinx a.) y=sin2x

OpenStudy (anonymous):

looks just like sine but period is \[\pi\] instead of \[2\pi\]

OpenStudy (anonymous):

why is that?

OpenStudy (anonymous):

because as x goes from 0 to pi, 2x goes from 0 to 2pi

OpenStudy (anonymous):

okay gotcha

OpenStudy (anonymous):

so what will the graph of sine look like?

OpenStudy (anonymous):

period of \[\sin(bx)\] is \[\frac{2\pi}{b}\]

OpenStudy (anonymous):

sorry, why?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=y%3Dsin%282x%29

OpenStudy (anonymous):

I don't think I understand periods

OpenStudy (anonymous):

same reason. sine is periodic with period 2pi as x goes from 0 to 2pi/b bx goes from 0 to 2pi

OpenStudy (anonymous):

period of sine is 2pi means \[\sin(x)=\sin(x+2\pi)\] you have just gone around the circle again

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

that is why graph goes up and down and up and down and up and down like it does

OpenStudy (anonymous):

i'm just so confused as to how I graph it knowing that it is periodic with 2π

OpenStudy (anonymous):

here is the graph of \[y=\sin(x)\] http://www.wolframalpha.com/input/?i=y%3Dsin%28x%29

OpenStudy (anonymous):

oh okay I got it because it crosses the x-axis everytime sine is 0 on the circle

OpenStudy (anonymous):

goes through (0,0) then up to 1, then down to 0, then down to -1, then back up to 0 at 2pi, then it repeats itself

OpenStudy (anonymous):

exactly. 0 at 0 then up to 1 at pi/2, then back to 0 at pi, then down to -1 at 3pi/2 then up to 0 at 2pi

OpenStudy (anonymous):

oh okay I got it because it crosses the x-axis everytime sine is 0 on the circle